# Insights from Large-Scale Communications: Initiation

### v1

@@ 4.1.2 Information Sharing at different redshifts
We use N-body simulations to study the information distribution at different redshifts in a $\Lambda CDM$ Universe. We went over $80$ N-body realisations of the distribution of DM particles at redshifts of 00, $0.5$, $1$, and $2$ ( see subsection 3.2 ). For a fair comparison, we compare the results with sample 3 (SDSS), since each cube has the same size of $250{\,h^{-1}\,{\rm Mpc}}$. Figure 8 shows that the overall amplitude of both information spectrum and mutual gain is much higher in sample 3 compared to the N-body cubes. One might expect the Sample 3 results to be substantially closer to $z=0.5$, but the mutual gain for Sample 3 is almost twice the $z=0.5$ distributions. The unfair comparison we are making between the information content of galaxy and DM distributions here is the cause of this difference. This is addressed in the next section. The characteristic that stands out in this case is the information content gradually increasing across all length scales with decreasing redshifts. Additionally, the homogeneity scale for the N-body cubes drops as we go towards higher redshifts. These findings are rather typical. The offset from a homogenous random background would increase with decreasing redshift, as the structures are better defined and less arbitrary on lower redshifts. Incoherent structures with greater uncertainty are less likely to contribute to the reduction of uncertainty at other distant points in the space. Hence, the scale of homogeneity is naturally expected to drop for distributions at higher redshifts. The results in Figure 8 are consistent with this expectation. For all the N-body cubes homogeneity is achieved within $130{\,h^{-1}\,{\rm Mpc}}$; whereas, the SDSS Sample 3 distribution tends towards homogeneity beyond $160{\,h^{-1}\,{\rm Mpc}}$.
@@ 4.1.3 Effect of linear bias and redshift-space distortions
Figure 9 presents the en bloc effect of RSD and bias on the information distribution. In the left panel of the figure, we show the information spectra of the distributions with and without RSD and biasing. Let us first compare the real-space and redshift-space unbiased distributions. We observe that the redshift-space distribution exhibits suppression of information on higher frequencies, whereas enhancement in the information spectrum can be noticed for larger wavelengths. The first effect is the manifestation of the Finger-of-God (FOG) effect, and the second one is caused by the Kaiser effect; both of these are effects of RSD. The random peculiar motion of galaxies in dense environments such as galaxy clusters causes the elongation of the apparent shape of structures along the line of sight. This FOG effect reduces the coherence of the particles in close vicinity, that exist in the real-space distribution, causing the signal to drop on higher frequencies. On the contrary, on large scales, galaxies experience bulk infall towards large overdensities, causing galaxies to appear closer to each other than they are. Hence artificial spatial coherence is introduced in the signal on large scales. This also explains the observed fall of information for Sample 1 in the top-left panel of Figure 7. Both the distance of the observer and the clustering of particles affect the degree of information suppression on large $k$. The suppression would increase with increasing density contrast and proximity of the observer relative to the cube. Unlike RSD, the effect of biasing does not affect the information in a high $k$ regime. Only the large-scale information gain is affected due to biasing. Information is found to be enhanced with biasing strength and this effect becomes more prominent as we move towards smaller $k$ values. However, the difference between the information spectrum of Sample 3 and the biased distribution for $k>0.1{\,h\,{\rm Mpc^{-1}}}$ is prominent and the difference increases with $k$. This tells us that a non-linear and scale-dependent biasing is required to model the information spectrum of the observed galaxy distributions, especially for a non-linear perturbation regime.
Figure 10: This figure shows the evolution of effective information entropy for a lognormal density distribution, from redshifts $20$ to the present day.
[…]
4.2 Evolution of Global Information entropy
In subsection 2.5 we discuss the method to measure the global information entropy at a given epoch for a log-normal density distribution and introduce three dimensionless quantities $\frac{d\ln{\cal G}^{U}}{d\ln a}$, $\frac{d\ln{\cal G}^{\prime}}{d\ln H}$ and $\frac{d\ln{\cal G}^{U}}{d\ln D_{+}}$ to analyse the evolution of effective information gain of the universe. These logarithmic derivatives measure how fast or slow the Information gain changes relative to the growth rate of matter density perturbation or expansion rate of the space. In Figure 10 we show these quantities as a function of redshift or scalefactor.
The top left panel of Figure 10 shows the evolution of the density distribution that causes the change in the global information gain. The middle-left panel shows the overall information entropy of the Universe at different epochs. The present Universe holds $\sim 2.8\times 10^{147}TB$, including the observed -ve information gain ( or information loss) due to structures formed at $z=0$. In the bottom-left panel, we show the rate of change of ${\cal I}$ as a function of the scalefactor. We notice that $\frac{{d\cal I}^{U}}{da}$ rapidly increases during the cosmic noon ($1.5\lesssim z\lesssim 3$) and then slows down to reach the maximum at $z=0.63$; This is the point at which the expansion of the universe switches from a decelerated to an accelerated phase. Between $z=2$ to $z=0.5$ information entropy undergoes a huge $\sim 300\%$ increment; whereas, it grows only $\sim 75\%$ from $z=0.5$ to $z=0$, as the matter domination gets over. At the current epoch, ${\cal I}^{U}$ is found to be increasing in a decelerated way. This is the opposite of what information gain is undergoing at present. The information loss appears to be accelerating as seen by the top-left panel. Although the information loss is significantly smaller than the effective information entropy till now, we find that $\sim 8\times 10^{144}TB$ information has already been used in large-scale structure formation. Note that this accounts only for the linear perturbations of the matter density field and the expenditure could be greater when one incorporates structures on smaller length scales ($<8{\,h^{-1}\,{\rm Mpc}}$). However, information gain is insensitive to perturbations beyond $\delta\sim 200$, so we expect the actual loss to be close to the measured value.
The second panel from the top shows the logarithmic derivative of information gain w.r.t the scale factor. The information gain increases $5$ times as fast as the scale factor in the early Universe but slows down with the end of matter domination. The Information gain is growing by $1.5$ times as fast as the expansion rate in present times. The second panel from the top shows how sensitive the relative change in information gain is w.r.t the acceleration of the Universe. Now $(-\,{\cal G}^{U})\leq 0$ and $H(a)\geq 0$ throughout, which leave us with $\frac{d(-{\cal G}^{U})}{dH}<0$. Hence, the information loss is overall decreasing with the growth in expansion rate and the relative change in the information loss is $2.5$ times faster than the relative change in the Hubble parameter. Moving to the bottom-right panel, we notice the relative growth of ${\cal G}^{U}$ to be $5$ times higher in comparison with the change in growing mode, at $z\geq 4$; it then gradually increased and becomes almost $9$ times of the relative change in the growing mode. The crucial thing to note here is $\frac{d(-{\cal G}^{U})}{dD_{+}}>0$. This suggests that despite the expansion of the Universe the growth of structures will always cause a loss in information unless $f(\Omega_{m})\leq 0$. Note that both information gain and information loss are associated with reduction of uncertainty. The -ve sign of information gain associates the effective entropic change with a higher degree of information reduction in underdense regions.
@@ 5 Conclusion
The two-point correlation function and its Fourier counterpart the power spectrum are the widely used statistics for studying galaxy clustering. However, these two estimators and their multiple variants that rely upon a finite number of moments of density perturbation, may not be sufficient in analysing the galaxy and dark-matter clustering in the non-linear regime. We introduce the information gain, an information-theoretic identity that uses all the higher-order moments of perturbation in analysing the nature of clustering in the observable galaxy distribution. We employ this tool to analyse the real and Fourier space distribution of information in 3 different galaxy samples prepared using data from SDSS, each representing a specific redshift zone and a particular class of galaxy. This provides an idea of the information flowing through different parts of the cosmic web. We also quantify the information reduction mutually caused by structures at any two points in space, in the accessible volume. We search for a universal scale of homogeneity, where ( by definition ) the mutual gain in information diminishes irrespective of the local clustering strength. The presence of such a characteristic scale is not found at least up to $130{\,h^{-1}\,{\rm Mpc}}$. Furthermore, to study the evolution of the Universe’s information content, we study a log-normal density distribution that evolves with time. With redshift-dependent variance, skewness and kurtosis, it mimics the probability distribution of the actual density contrast in the matter density field. It is found that the growth of structures inevitably causes a reduction in information, irrespective of the expansion rate of the Universe. Although we have limited our discussion to the $\Lambda CDM$ cosmology, it is possible to examine the global information content of different cosmological models and constrain the cosmological parameters with the aid of appropriate observables. In this work, we have assumed that the characteristic change in total information entropy after $z=0.6$ is due to the matter component being dominated by dark energy. However, a converse scenario might also be true. Entropy being one of the fundamental physical entities in the Universe could cause the observed accelerated expansion itself. Detailed studies are required to comprehend the actual reason. This paper is meant to be the initial step, laying the foundation for future research that will build on it and will be better equipped to address these issues.
@@ 7 Acknowledgement
I thank Prof. Somnath Bharadwaj, Prof. Sayan Kar, Dr. Biswajit Pandey and fellow researchers at IIT-Kharagpur, for their suggestions and useful discussions. Special thanks to DST-SERB for support through the National Post-Doctoral Fellowship (PDF/2022/000149). Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingrand-designom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
[…]
Appendix A Shot noise correction for Information gain
@@
Consider a discrete distribution of ${\cal N}$ galaxies confined in a finite cubic volume, where each of the voxels accommodates $\bar{f}_{\cal M}\,\cal{N}$ galaxies on average. From Equation 15 for any $i^{th}$ voxel we get the information gain as
${\cal G}^{(i)}=\frac{K}{U_{0}\bar{I}}\left[\Delta I^{(i)}\,\exp\{\,-\Delta I^{(i)}\,\}\right]$ (36)
@@
$\Delta I_{SNC}(\delta)=\Delta I(\delta)-\left|\frac{\partial\,\Delta I}{\partial n}\right|\sigma_{n}$ (37)
Now, $\frac{\partial\,\Delta I}{\partial n}=\frac{1}{\delta_{c}\,\bar{f}_{\cal M}\,\cal{N}}$, and if we assume the particle count to follow a Poisson distribution, then $\sigma_{n}=\sqrt{\bar{f}_{\cal M}\,\cal{N}}$. Hence, the shot noise correction would lead us to
${\cal G}^{(i)}_{SNC}(\eta)$ $=$ $\frac{K}{\bar{I}\,U_{0}^{SNC}}\left[\Delta I^{(i)}_{SNC}\,\exp\{\,-\Delta I^{(i)}_{SNC}\,\}\right]$ (38)
@@
$\Delta I^{(i)}_{SNC}=\Delta I^{(i)}-\frac{1}{\delta_{c}\sqrt{\bar{f}_{\cal M}\,\cal{N}}}$ (39)
Appendix B Finding $n^{th}$ order derivative of $k^{th}$ EWM
Let us start by writing the $k^{th}$ EWM as
${\cal A}_{k}=\frac{U_{k}}{U_{0}}$ (40)
Where
$U_{k}=\int_{\mathbb{R}^{3}}\eta^{k}e^{-\eta}d^{3}x$ (41)
By definition $\eta(\mathbf{x},a)=D_{+}(a)\frac{\delta(\mathbf{x})}{\delta_{c}}$; hence we can deduce
$\mathbb{D}^{q}\eta=\eta Q_{q}.$ (42)
by denoting $\,\,\frac{d^{q}}{da^{q}}$ as $\mathbb{D}^{q}\,\,$ and $\,\,\frac{1}{D_{+}}\frac{\partial^{q}D_{+}}{\partial a^{q}}$ as $Q_{q}\,\,$. Note that $D_{+}$ being a function of $a$ alone allows the partial and total derivatives to have interchangeable connotations. Equation 42 also leads to the identities
$Q_{0}$ $=$ $1,$
$Q_{1}$ $=$ $\frac{f(\Omega_{m})}{a},$
$\mathbb{D}^{1}Q_{n}$ $=$ $Q_{n+1}-Q_{n}Q_{1},$
$\mathbb{\mathbb{missing}}{D}^{q}Q_{n}$ $=$ ${D}^{q-1}Q_{n+1}-\sum_{j=0}^{q-1}{}^{q-1}C_{j}\,\,\,\mathbb{D}^{q-j-1}Q_{n}\,\,\,\mathbb{D}^{j-1}Q_{1}.$
We use Leibniz’s general product rule of differentiation to get the last equation.
Using Equation 43 the first derivative of $U_{k}$ is found as
$\mathbb{D}^{1}U_{k}=Q_{1}\,[\,kU_{k}-U_{k+1}\,],$ (43)
and the $1^{st}$ derivative of ${\cal A}_{k}$ would be
$\mathbb{D}^{1}{\cal A}_{k}$ $=$ $\left[\,\frac{\mathbb{D}^{1}U_{k}}{U_{0}}-\frac{U_{k}\,\mathbb{D}^{1}U_{0}}{U^{2}_{0}}\,\right]$ (44)
$=$ $Q_{1}\,[\,k{\cal A}_{k}-{\cal A}_{k+1}+{\cal A}_{1}{\cal A}_{k}\,]$ (45)
Any $n^{th}$ order derivative of ${\cal A}_{k}$ can now be written as
$\mathbb{D}^{n}{\cal A}_{k}=\mathbb{D}^{n-1}\left\{Q_{1}\,[\,k{\cal A}_{k}-{\cal A}_{k+1}+{\cal A}_{1}{\cal A}_{k}\,]\right\}$ (46)
Finally, the $n^{th}$ derivative ( w.r.t $a$) of the $k^{th}$ EWM is found in form of the general recurrence relation
$\mathbb{D}^{n}{\cal A}_{k}$ $=$ $\sum_{q=0}^{n-1}\,\,{}^{n-1}C_{q}\,\,\,\mathbb{F}^{n-q-1}_{1}\cdot\mathbb{T}^{q}_{k}.$ (47)
$\text{where}\,\,\,\mathbb{F}^{n-q-1}_{s}$ $=$ $\mathbb{D}^{n-q-1}Q_{s}\,\,$
One can find the following identities using Equation 47.
$\mathbb{D}^{1}{\cal A}_{1}$ $=$ $Q_{1}\left\{{\cal A}_{1}-{\cal A}_{2}+{\cal A}^{2}_{1}\right\}$
$\mathbb{D}^{1}{\cal A}_{2}$ $=$ $Q_{1}\left\{2{\cal A}_{2}-{\cal A}_{3}+{\cal A}_{2}{\cal A}_{1}\right\}$
$\mathbb{D}^{2}{\cal A}_{1}$ $=$ $Q_{1}^{2}\left\{2{\cal A}_{1}^{3}+{\cal A}_{1}^{2}-3{\cal A}_{1}{\cal A}_{2}-2{\cal A}_{2}+{\cal A}_{3}\right\}$
$+Q_{2}\left\{{\cal A}_{1}-{\cal A}_{2}+{\cal A}^{2}_{1}\right\}$
Appendix C Scale independent estimation of linear bias

### v2

@@ 4.2.2 Information Sharing at different redshifts
We use N-body simulations to study the information distribution at different redshifts in a $\Lambda CDM$ Universe. We went over $80$ N-body realisations of the distribution of DM particles at redshifts of 00, $0.5$, $1$, and $2$ ( see subsection 3.2 ). For a fair comparison, we compare the results with sample 3 (SDSS), since each cube has the same size of $250{\,h^{-1}\,{\rm Mpc}}$. Figure 9 shows that the overall amplitude of both information spectrum and mutual gain is much higher in sample 3 compared to the N-body cubes. One might expect the Sample 3 results to be substantially closer to $z=0.5$, but the mutual gain for Sample 3 is almost twice the $z=0.5$ distributions. The cause of this difference is the unfair comparison between galaxy and DM distributions. One has to make a fair comparison by using a mock galaxy distribution instead; this is addressed in the next section. The characteristic that stands out in this case is the information content gradually increasing across all length scales with decreasing redshifts. Additionally, the homogeneity scale for the N-body cubes drops as we go towards higher redshifts. These findings are rather typical. The offset from a homogenous random background would increase with decreasing redshift, as the structures are better defined and less arbitrary on lower redshifts. Incoherent structures with greater uncertainty are less likely to contribute to the reduction of uncertainty at other distant points in the space. Hence, the scale of homogeneity is naturally expected to drop for distributions at higher redshifts. The results in Figure 9 are consistent with this expectation. For all the N-body cubes homogeneity is achieved within $130{\,h^{-1}\,{\rm Mpc}}$; whereas, the SDSS Sample 3 distribution tends towards homogeneity beyond $150{\,h^{-1}\,{\rm Mpc}}$. An oscillating residue is noticed beyond that which may or may not originate due to the finite volume of the cube.
@@ 4.2.3 Effect of linear bias and redshift-space distortions
Figure 10 presents the en bloc effect of RSD and bias on the information distribution. In the left panel of the figure, we show the information spectra of the distributions with and without RSD and biasing. Let us first compare the real-space and redshift-space unbiased distributions. We observe that the redshift-space distribution exhibits suppression of information on higher frequencies, whereas enhancement in the information spectrum can be noticed for larger wavelengths. The first effect is the manifestation of the Finger-of-God (FOG) effect, and the second one is caused by the Kaiser effect; both of these are effects of RSD. The random peculiar motion of galaxies in dense environments such as galaxy clusters causes the elongation of the apparent shape of structures along the line of sight. This FOG effect reduces the coherence of the particles in close vicinity, that exist in the real-space distribution, causing the signal to drop on higher frequencies. On the contrary, on large scales, galaxies experience bulk infall towards large overdensities, causing galaxies to appear closer to each other than they are. Hence artificial spatial coherence is introduced in the signal on large scales. This also explains the observed fall of information for Sample 1 in the top-left panel of Figure 8. Both the distance of the observer and the clustering of particles affect the degree of information suppression on large $k$. The suppression would increase with increasing density contrast and proximity of the observer relative to the cube. Unlike RSD, the effect of biasing does not affect the information in a high $k$ regime. Only the large-scale information gain is affected due to biasing. Information is found to be enhanced with biasing strength and this effect becomes more prominent as we move towards smaller $k$ values. However, the difference between the information spectrum of Sample 3 and the biased distribution for $k>0.1{\,h\,{\rm Mpc^{-1}}}$ is prominent and the difference increases with $k$. This indicates the necessity of non-linear and scale-dependent biasing to model the information spectrum of the observed galaxy distributions, especially for a non-linear perturbation regime.
4.2.4 On the scale of homogeneity
On the right panel of this figure, we show the information sharing for the different distributions. We also notice the enhancement in uncertainty when anti-biasing is applied, this reduces the entropic gain for the distribution with $b_{1}=0.75$, in all wavelengths. As we keep the bias increasing, the dark-matter particles are stripped away from the existing uncertainty, and a significant entropic gain is observed. For $b_{1}=2$, we find the mutual gain of the redshifted distribution to come within 1-$\sigma$ of the SDSS Sample 3. Despite the differences on the largest scales, a reduced-$\chi^{2}$ of $0.49$ suggests that the two distributions are drawn from the same population, with a $99.8\%$ confidence level. Both the real-space unbiased and redshift-space anti-biased distributions reach homogeneity within $120{\,h^{-1}\,{\rm Mpc}}$ whereas the distributions with $b_{1}=1$, $b_{1}=1.5$ and $b_{1}=2$ in redshift space achieve homogeneity around $150-160{\,h^{-1}\,{\rm Mpc}}$. Compared with the $b_{1}=2$ distribution one can expect the SDSS sample 3 distribution to reach homogeneity beyond $160{\,h^{-1}\,{\rm Mpc}}$. Although large standard deviations beyond $100{\,h^{-1}\,{\rm Mpc}}$ do not allow us to have a unanimous confirmation on a global homogeneity scale within this scope, we can at least rule out its presence in the observable Universe up to $130{\,h^{-1}\,{\rm Mpc}}$.
@@ 5 Conclusion
The two-point correlation function and its Fourier counterpart the power spectrum are the widely used statistics for studying galaxy clustering. However, these two estimators and their multiple variants that rely upon a finite number of moments of density perturbation, may not be sufficient in analysing the galaxy and dark-matter clustering in the non-linear regime. We introduce the entropic gain, an information-theoretic identity that uses all the higher-order moments of perturbation in analysing the nature of clustering in the observable galaxy distribution. We employ this tool to analyse the real and Fourier space distribution of information in 3 different galaxy samples prepared using data from SDSS, each representing a specific redshift zone and a particular class of galaxy. We also quantify the information reduction mutually caused by structures at any two points in space, in the accessible volume. We search for a universal scale of homogeneity, where ( by definition ) the mutual gain in information diminishes irrespective of the local clustering strength. We don’t find such a characteristic scale at least up to $130{\,h^{-1}\,{\rm Mpc}}$. Furthermore, to study the evolution of the Universe’s information content, we study a log-normal density distribution that evolves with time. Redshift-dependent variance, skewness and kurtosis mimic the probability distribution of the actual density contrast in the matter density field. It is found that the growth of structures inevitably causes a reduction in information entropy. Entropy being one of the fundamental physical entities in the Universe could cause the observed accelerated expansion itself. For a thermodynamic system, the increment in the volume allows it to have a larger number of microstates to access. A configuration with a higher volume is preferred by a system due to its innate inclination towards achieving a higher entropy state. The nonlinear reduction in entropy due to structure formation can indirectly cause the accelerated expansion of the space. However, here we are measuring merely the drop in information entropy. So the theory of the accelerated expansion being caused by a growing entropic drop due to structure formation requires a deeper inspection. Although we have limited our discussion to the $\Lambda CDM$ cosmology, it is possible to study the entropic gain of different cosmological models and constrain the cosmological parameters with the aid of appropriate observables. This paper is meant to be the initial step, laying the foundation for future research that will build on it and will be better equipped to address these issues.
@@ 7 Acknowledgement
I thank Prof. Somnath Bharadwaj, Prof. Sayan Kar, Dr. Biswajit Pandey, Dr. Maret Einasto and fellow researchers at IIT-Kharagpur, for their suggestions and useful discussions. Special thanks to DST-SERB for support through the National Post-Doctoral Fellowship (PDF/2022/000149). Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingrand-designom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
Appendix A Appendix:
A.1 Finding $n^{th}$ order derivative of $k^{th}$ EWM
Let us start by writing the $k^{th}$ EWM as
${\cal A}_{k}=\frac{U_{k}}{U_{0}}$ (A.1)
Where
$U_{k}=\int_{\mathbb{R}^{3}}\eta^{k}e^{-\eta}d^{3}x$ (A.2)
By definition $\eta(\mathbf{x},a)=D_{+}(a)\frac{\delta(\mathbf{x})}{\delta_{c}}$; hence by denoting $\,\,\frac{d^{q}}{da^{q}}$ as $\mathbb{D}^{q}\,\,$ and $\,\,\frac{1}{D_{+}}\frac{\partial^{q}D_{+}}{\partial a^{q}}$ as $Q_{q}\,\,$ we can write
$\mathbb{D}^{q}\eta=\eta Q_{q}.$ (A.3)
Note that $D_{+}$ being a function of $a$ alone allows the partial and total derivatives to have interchangeable connotations. Equation A.3 also leads to the identities
$Q_{0}$ $=$ $1,$
$Q_{1}$ $=$ $\frac{f(\Omega_{m})}{a},$
$Q_{n+1}$ $=$ $\mathbb{D}^{1}Q_{n}+Q_{n}Q_{1},$
Using Equation A.2 and Equation A.3 the first derivative of $U_{k}$ is found as
$\mathbb{D}^{1}U_{k}=Q_{1}\,[\,kU_{k}-U_{k+1}\,].$ (A.4)
The $1^{st}$ derivative of ${\cal A}_{k}$ then would be
$\mathbb{D}^{1}{\cal A}_{k}$ $=$ $\left[\,\frac{\mathbb{D}^{1}U_{k}}{U_{0}}-\frac{U_{k}\,\mathbb{D}^{1}U_{0}}{U^{2}_{0}}\,\right]$ (A.5)
$=$ $Q_{1}\,[\,k{\cal A}_{k}-{\cal A}_{k+1}+{\cal A}_{1}{\cal A}_{k}\,]$
Any the $n^{th}$ order derivative of the $k^{th}$ EWM can now be estimated from the general recurrence relation
$\mathbb{D}^{n}{\cal A}_{k}$ $=$ $\sum_{q=0}^{n-1}\,\,{}^{n-1}C_{q}\,\,\,\mathbb{F}^{n-q-1}_{1}\cdot\mathbb{T}^{q}_{k}.$ (A.6)
$\text{where}\,\,\,\mathbb{F}^{n-q-1}_{s}$ $=$ $\mathbb{D}^{n-q-1}Q_{s}\,\,$
One can find the following identities using Equation A.6.
$\mathbb{D}^{1}{\cal A}_{1}$ $=$ $Q_{1}\left\{{\cal A}_{1}-{\cal A}_{2}+{\cal A}^{2}_{1}\right\}$ (A.7)
$\mathbb{D}^{1}{\cal A}_{2}$ $=$ $Q_{1}\left\{2{\cal A}_{2}-{\cal A}_{3}+{\cal A}_{2}{\cal A}_{1}\right\}$ (A.8)
$\mathbb{D}^{2}{\cal A}_{1}$ $=$ $Q_{1}^{2}\left\{2{\cal A}_{1}^{3}+{\cal A}_{1}^{2}-3{\cal A}_{1}{\cal A}_{2}-2{\cal A}_{2}+{\cal A}_{3}\right\}+Q_{2}\left\{{\cal A}_{1}-{\cal A}_{2}+{\cal A}^{2}_{1}\right\}$ (A.9)
A.2 Shot noise correction for entropic gain
@@
Consider a discrete distribution of ${\cal N}$ galaxies confined in a finite cubic volume, where each of the voxels accommodates $\bar{f}_{\cal M}\,\cal{N}$ galaxies on average. From Equation 2.13 for any $i^{th}$ voxel we get the specific information gain as
$\Phi\,(\,\mathbf{x}_{i}\,)$ $=$ $\frac{1}{\bar{I}\,U_{0}\,\Delta x^{3}}\biggl{[}\,\,\,\Delta I\,(\delta\,(\mathbf{x}_{i})\,)\,\,\exp\{\,-\Delta I\,(\,\delta(\mathbf{x}_{i})\,)\,\}\,\,\,\biggr{]}.$ (A.10)
@@
$\Delta I_{SNC}(\delta)=\Delta I(\delta)-\left|\frac{\partial\,\Delta I}{\partial n}\right|\sigma_{n}$ (A.11)
We also have $\frac{\partial\,\Delta I}{\partial n}=\frac{1}{\delta_{c}\,\bar{f}_{\cal M}\,\cal{N}}$ and $\sigma_{n}=\sqrt{\bar{f}_{\cal M}\,\cal{N}}$ assumimng the particle count follows a Poisson distribution. Hence, the shot noise correction would give us
$\Phi^{(i)}_{SNC}(\eta)$ $=$ $\frac{1}{\bar{I}\,U_{0}^{SNC}\Delta x^{3}}\left[\Delta I^{(i)}_{SNC}\,\exp\{\,-\Delta I^{(i)}_{SNC}\,\}\right]$ (A.12)
@@
$\Delta I^{(i)}_{SNC}=\Delta I^{(i)}-\frac{1}{\delta_{c}\sqrt{\bar{f}_{\cal M}\,\cal{N}}}$ (A.13)
A.3 Scale independent estimation of linear bias